OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7 Would it be -14?

OpenStudy (anonymous):

@jigglypuff314?

OpenStudy (jdoe0001):

A \(\bf \color{brown}{ number}\) divided by 7 is 6 less than –8.\\ \(\bf \cfrac{{\color{brown}{ x}}}{7}=-8-6\)

OpenStudy (domebotnos):

6 less than -8 is -14. A shortcut would be: -14*7=-98 To check the answer do -98/7=-14 The answer is not x = -14.

OpenStudy (jdoe0001):

solve for "x"

OpenStudy (anonymous):

wo wo wo, i might be wrong then

OpenStudy (domebotnos):

A. -98 is the only possible correct answer.

OpenStudy (anonymous):

yeah i'm bad, sorry

OpenStudy (anonymous):

Oh thanks

OpenStudy (domebotnos):

Np. :)

OpenStudy (anonymous):

I need help with this one too Which line shows the first error in the solution? 6x+4=3x-8 6x-3x+4=3x-3x-8 3x+4=-8 3x+4-4=-8-4 3x=-4 3x/3=-4/3 x=-4 A Line 3 B Line 4 C Line 5 D Line 6

OpenStudy (anonymous):

The best method, is to solve it by yourself, side by side with that work that you just posted, and see when you notice that you got a different result

OpenStudy (domebotnos):

Look closely at line 5. -8-4 does not equal -4, it equals -12. C. Line 5.